OpenStudy (teacherspet02):

what is the area of a regular pentagon, with an apothem 14 in long and a side length 20 in long? a. 1400 in^2 b.700 in^2 c.175 in^2 d.140 in^2

OpenStudy (teacherspet02):

@Will.H

OpenStudy (will.h):

Regular pentagon are\[A=1/4 \sqrt{5(5+2\sqrt{5)}a^2}\] Therefore after plugging a= 20 Are would equal 688.19

OpenStudy (michele_laino):

hint: we can apply this formula: \[area = \frac{{erimeter \times apothem}}{2}\]

OpenStudy (michele_laino):

oops.. \[area = \frac{{perimeter \times apothem}}{2}\]

OpenStudy (teacherspet02):

i got tht when i did it to, but i dont know what to do next bc that isnt an answer they want....?

OpenStudy (michele_laino):

wherein: \(perimeter=20 \times 5=...\)

OpenStudy (teacherspet02):

So A?

OpenStudy (will.h):

@Michele_Laino i feel like am missing something, what is it?

OpenStudy (teacherspet02):

ohh i forgot to divide... thank you!

OpenStudy (will.h):

Good job

OpenStudy (michele_laino):

please wait

OpenStudy (michele_laino):

I have made a typo: \[area = \frac{{perimeter \times apothem}}{2} = \frac{{100 \times 14}}{2} = 700\]

OpenStudy (teacherspet02):

700 is right?

OpenStudy (michele_laino):

that's right!

OpenStudy (will.h):

Nicely done Michele

OpenStudy (michele_laino):

thanks!! :) @Will.H